# xLSTM-Mixer: Multivariate Time Series Forecasting by Mixing via Scalar Memories

### v2

@@ Abstract
Time series data is prevalent across numerous fields, necessitating the development of robust and accurate forecasting models. Capturing patterns both within and between temporal and multivariate components is crucial for reliable predictions. We introduce xLSTM-Mixer, a model designed to effectively integrate temporal sequences, joint time-variate information, and multiple perspectives for robust forecasting. Our approach begins with a linear forecast shared across variates, which is then refined by xLSTM blocks. They serve as key elements for modeling the complex dynamics of challenging time series data. xLSTM-Mixer ultimately reconciles two distinct views to produce the final forecast. Our extensive evaluations demonstrate its superior long-term forecasting performance compared to recent state-of-the-art methods. A thorough model analysis provides further insights into its key components and confirms its robustness and effectiveness. This work contributes to the resurgence of recurrent models in time series forecasting.
@@ 1 Introduction
Time series are an essential data modality ubiquitous in many critical fields of application, such as medicine (Hosseini et al., 2021), manufacturing (Essien & Giannetti, 2020), logistics (Seyedan & Mafakheri, 2020), traffic management (Lippi et al., 2013), finance (Lin et al., 2012), audio processing (Latif et al., 2023), and weather modeling (Lam et al., 2023). While significant progress in time series forecasting has been made over the decades, the field is still far from being solved. The regular appearance of yet better models and improved combinations of existing approaches exemplifies this. Further increasing the forecast quality obtained from machine learning models promises a manifold of improvements, such as higher efficiency in manufacturing and transportation as well as more accurate medical treatments.
@@
Historically, recurrent neural networks (RNNs) and their powerful successors were natural choices for deep learning-based time series forecasting (Hochreiter & Schmidhuber, 1997; Cho et al., 2014). Today, large Transformers (Vaswani et al., 2017) are applied extensively to time series tasks, including forecasting. Many improvements to the vanilla architecture have since been proposed, including patching (Nie et al., 2023), decompositions (Zeng et al., 2023), and tokenization inversions (Liu et al., 2023). However, some of their limitations are yet to be lifted. For instance, they typically require large datasets to train successfully, restricting their use to only a subset of conceivable applications. Furthermore, they are inefficient when applied to long sequences due to the cost of the attention mechanism being quadratic in the number of variates and time steps, depending on the specific choice of tokenization. Therefore, recurrent and state space models (SSMs) (Patro & Agneeswaran, 2024) are experiencing a resurgence of interest in overcoming such limitations. Specifically, Beck et al. (2024) revisited recurrent models by borrowing insights gained from Transformers applied to many domains, specifically to natural language processing. They propose Extended Long Short-Term Memory (xLSTM) models as a viable alternative to current sequence models.
@@
We propose xLSTM-Mixer [^1] https://github.com/mauricekraus/xLSTM-Mixer, a new state-of-the-art method for time series forecasting using recurrent deep learning methods. Specifically, we augment the highly expressive xLSTM architecture with carefully crafted time, variate, and multi-view mixing. These operations regularize the training and limit the model parameters by weight-sharing, effectively improving the learning of features necessary for accurate forecasting. xLSTM-Mixer initially computes a channel-independent linear forecast shared over the variates. It is then up-projected to a higher hidden dimension and subsequently refined by an xLSTM stack. It performs multi-view forecasting by producing a forecast from the original and reversed up-projected embedding. The powerful xLSTM cells thereby jointly mix time and variate information to capture complex patterns from the data. Both forecasts are eventually reconciled by a learned linear projection into the final prediction, again by mixing time. An overview of our method is shown in Figure 1.
@@
Figure 1: The xLSTM-Mixer architecture consists of three stages: (1) An initial NLinear forecast assuming channel independence and performing time mixing; (2) subsequent joint mixing, which mixes variate and time information through crucial applications of sLSTM blocks in two views; and (3) a final view mixing, where the two latent forecast views are reconciled into a coherent final forecast.
@@
    We propose xLSTM-Mixer, a state-of-the-art method for time series forecasting using recurrent deep learning methods.
@@
    We extensively compare xLSTM-Mixer with existing methods for multivariate long-term time series forecasting and perform in-depth model analyses. The experiments demonstrate that xLSTM-Mixer consistently achieves state-of-the-art performance in a wide range of benchmarks.
The following work is structured as follows: In the upcoming Sec. 2, we introduce preliminaries to then motivate and explain xLSTM-Mixer in Sec. 3. We then present comprehensive experiments on its effectiveness and inner workings in Sec. 4. We finally contextualize the findings within the related work in Sec. 5 and close with a conclusion and outlook in Sec. 6.
@@ 2 Background
After introducing the notation used throughout this work, we review xLSTM blocks and discuss leveraging channel mixing or their independence in time series models.
2.1 Notation
In multivariate time series forecasting, the model is presented with a time series $\bm{X}=\left(\bm{x}_{1},\dots,\bm{x}_{T}\right)\in\mathbb{R}^{V\times T}$ consisting of $T$ time steps with $V$ variates each. Given this context, the forecaster shall predict the future values $\bm{Y}=\left(\bm{x}_{T+1},\dots,\bm{x}_{T+H}\right)\in\mathbb{R}^{V\times H}$ up to a horizon $H$. A variate (also called a channel) can be any scalar measurement, such as the occupancy of a road or the oil temperature in a power plant. The measurements are assumed to be carried out jointly, such that the $T+H$ time steps reflect a regularly sampled multivariate signal. A time series dataset consists of $N$ such pairs $\left\{\left(\bm{X}^{(i)},\bm{Y}^{(i)}\right)\right\}_{i\in\{1,\dots,N\}}$ divided into train, validation, and test portions.
@@ 2.2 Extended Long Short-Term Memory (xLSTM)
Beck et al. (2024) propose xLSTM architectures consisting of two building blocks, namely the sLSTM and mLSTM modules. To harness the full expressivity of xLSTMs within each step and across the computation sequence, we employ a stack of sLSTM blocks without any mLSTM blocks. The latter are less suited for joint mixing due to their independent treatment of the sequence elements, making it impossible to learn any relationships between them directly. We will continue by recalling the construction of sLSTM cells.
The standard LSTM architecture of Hochreiter & Schmidhuber (1997) involves updating the cell state $\mathbf{c}_{t}$ through a combination of input, forget, and output gates, which regulate the flow of information across tokens. sLSTM blocks enhance this by incorporating exponential gating and memory mixing (Greff et al., 2017) to handle complex temporal and cross-variate dependencies more effectively. The sLSTM updates the cell $\bm{c}_{t}$ and hidden state $\bm{h}_{t}$ using three gates as follows:
$\bm{c}_{t}$ $=\bm{f}_{t}\odot c_{t-1}+\bm{i}_{t}\odot\bm{z}_{t}$ cell state (1)
@@
$\bm{z}_{t}$ $=\tanh\bigl{(}\bm{W}_{z}\bm{x}_{t}+\bm{R}_{z}h_{t-1}+\bm{b}_{z}\bigr{)}$ cell input (4)
@@
In this setup, the matrices $\bm{W}_{z},\bm{W}_{i},\bm{W}_{f},$ and $\bm{W}_{o}$ are input weights mapping the input token $\bm{x}_{t}$ to the cell input $\bm{z}_{t}$, input gate, forget gate, and output gate, respectively. The states $\bm{n}_{t}$ and $\bm{m}_{t}$ serve as necessary normalization and training stabilization, respectively.
As Beck et al. have shown, it is beneficial to restrict the memory mixing performed by the recurrent weight matrices $\bm{R}_{z},\bm{R}_{i},\bm{R}_{f},$ and $\bm{R}_{o}$ to individual heads, inspired by the multi-head setup of Transformers (Zeng et al., 2023), yet more restricted and therefore more efficient to compute. In particular, each token gets broken up into separate pieces, where the input weights $\bm{W}_{z,i,f,o}$ act across all of them, but the recurrence matrices $\bm{R}_{z,i,f,o}$ are implemented as block-diagonals and therefore only act within each piece. This permits specialization of the individual heads to patterns specific to the respective section of the tokens and empirically does not sacrifice expressivity.
2.3 Channel Independence and Mixing in Time Series Models
Multiple works have investigated whether it is beneficial to learn representations of the time and variate dimensions jointly or separately. Intuitively, because joint mixing is strictly more expressive, one might think it should always be preferred. It is indeed used by many methods such as Temporal Convolutional Networks (TCN) (Lea et al., 2016), N-BEATS (Oreshkin et al., 2019), N-HiTS (Challu et al., 2023), and many Transformers (Vaswani et al., 2017), including Temporal Fusion Transformer (TFT) (Lim et al., 2021), Autoformer (Wu et al., 2021), and FEDFormer (Zhou et al., 2022). However, treating slices of the input data independently assumes an invariance to temporal or variate positions and serves as a strong regularization against overfitting, reminiscent of kernels in CNNs. Prominent models implementing some aspects of channel independence in multivariate time series forecasting are PatchTST (Nie et al., 2023) and iTransformer (Liu et al., 2023). TiDE (Das et al., 2023), on the other hand, contains a time-step shared feature projection and temporal decoder but treats variates jointly. As Tolstikhin et al. (2021) have shown with MLP-Mixer, interleaving mixing of all channels in each token and all tokens per channel does not empirically sacrifice any expressivity and instead improves performance. This idea has since been applied to time series, too, namely in architectures such as TimeMixer (Chen et al., 2023c) and TSMixer (Chen et al., 2023c), and is one of the key components of our method xLSTM-Mixer.
@@ 2.3 Channel Independence and Mixing in Time Series Models
Now we have everything at hand to introduce xLSTM-Mixer as depicted in Fig. 1. It carefully integrates three key components: (1) an initial linear forecast with time mixing, (2) joint mixing using powerful sLSTM modules, and (3) an eventual combination of two views by a final fully connected layer. The transposing steps between the key components enable capturing complex temporal and intra-variate patterns while facilitating easy trainability and limiting parameter counts. The sLSTM block, in particular, can learn intricate non-linear relationships hidden within the data along both the time and variate dimensions. The xLSTM-Mixer architecture is furthermore equipped with normalization layers and skip connections to improve training stability and overall effectiveness.
3.1 Key Component 1: Normalization and Initial Linear Forecast
Normalization has become an essential ingredient of modern deep learning architectures (Huang et al., 2023). For time series in particular, reversible instance norm (RevIN) (Kim et al., 2022) is a general recipe for improving forecasting performance, where each time series instance is normalized by its mean and variance and furthermore scaled and offset by learnable scalars $\gamma$ and $\beta$:
$\bm{x}_{t}^{\text{norm}}=\operatorname{RevIN}(\bm{x}_{t})=\gamma\left(\frac{\bm{x}_{t}-\mathbb{E}\left[\bm{x}\right]}{\sqrt{\mathrm{Var}\left[\bm{x}\right]}+\epsilon}\right)+\beta.$
We apply it as part of xLSTM-Mixer, and at the end of the entire pipeline, we invert the RevIN operation to obtain the final prediction. In the case of xLSTM-Mixer, the typical skip connections found in mixer acrchitectures (Tolstikhin et al., 2021; Chen et al., 2023c) are taken up by RevIN, the normalization in the NLinear forecast explained shortly, and the integral skip connections within each sLSTM block.
@@ 3.1 Key Component 1: Normalization and Initial Linear Forecast
where $\operatorname{FC}(\cdot)$ denotes a fully-connected linear layer with bias term. Sharing this model across variates limits parameter counts, and the weight-tying serves as a useful regularization. The quality of this initial forecast will be investigated in Sec. 4.1 and 4.2.
3.2 Key Component 2: sLSTM Refinement
@@
While the NLinear forecast $\bm{x}^{\text{initial}}$ captures the basic patterns between the historic and future time steps, its quality alone is insufficient for today’s challenging time series datasets. We, therefore, refine it using powerful sLSTM blocks. As a first step, it is crucial to increase the embedding dimension of the data to provide enough latent dimensions $D$ for the sLSTM cells: $\bm{x}^{\text{up}}=\operatorname{FC}^{\text{up}}\left(\bm{x}^{\text{initial}}\right)$. This pre-up-projection is similar to what is commonly performed in SSMs (Beck et al., 2024). We weight-share $\operatorname{FC}^{\text{up}}$ across variates to perform time-mixing similar to the initial forecast. Note that this step does not maintain the temporal ordering within the embedding token dimensions, as was the case up until this step, and instead embeds it into a higher latent dimension.
@@
The stack of $M$ sLSTM blocks $\mathcal{S}(\cdot)$ transforms $\bm{x}^{\text{up}}$ as defined in Eq. 1 to 8. The recurrent model strides over the data in variate order, i.e., where each token represents all time steps from a single variate as in the work of Liu et al. (2023). The sLSTM blocks learn intricate non-linear relationships hidden within the data along both the time and variate dimensions. The mixing of the hidden state is still limited to blocks of consecutive dimensions, aiding efficient learning and inference while allowing for effective cross-variate interaction during the recurrent processing. Striding over variates has the benefit of linear runtime scaling in the number of variates at a constant number of parameters. It, however, comes at the cost of possibly fixing a suboptimal order of variates. While this is empirically not a significant limitation, we leave investigations into how to find a suitable ordering for future work. In addition to a large embedding dim, we observed a high number of heads being crucial for effective forecasting.
@@
The sLSTM cells’ first hidden state $\bm{h}_{t-1}$ must be initialized before each sequence of tokens can be processed. Extending the initial description of these blocks, we propose learning a single initial embedding token $\bm{\eta}\in\mathbb{R}^{D}$ that gets prepended to each encoded time series $\bm{x}^{\text{up}}$. These initial embeddings draw from recent advances in Large Language Models, where learnable "soft prompt" tokens are used to condition models and improve their ability to generate coherent outputs (Lester et al., 2021; Li & Liang, 2021; Chen et al., 2023a; b). Recent research has extended the application of soft prompts to LLM-based time series forecasting (Cao et al., 2023; Sun et al., 2024), emphasizing their adaptability and effectiveness in improving model performance across modalities. These tokens enable greater flexibility and conditioning, allowing the model to adapt its initial memory representation to specific dataset characteristics and to dynamically interact with the time and variate data. Soft prompts can be readily optimized through back-propagation with very little overhead.
3.3 Key Component 3: Multi-View Mixing

### v3

@@ Abstract
Time series data is prevalent across numerous fields, necessitating the development of robust and accurate forecasting models. Capturing patterns both within and between temporal and multivariate components is crucial for reliable predictions. We introduce xLSTM-Mixer, a model designed to effectively integrate temporal sequences, joint time-variate information, and multiple perspectives for robust forecasting. Our approach begins with a linear forecast shared across variates, which is then refined by xLSTM blocks. They serve as key elements for modeling the complex dynamics of challenging time series data. xLSTM-Mixer ultimately reconciles two distinct views to produce the final forecast. Our extensive evaluations demonstrate its superior long-term forecasting performance compared to recent state-of-the-art methods while requiring very little memory. A thorough model analysis provides further insights into its key components and confirms its robustness and effectiveness. This work contributes to the resurgence of recurrent models in forecasting by combining them, for the first time, with mixing architectures.
@@ 1 Introduction
Figure 1: xLSTM-Mixer provides excellent forecasts with a very low memory footprint while being sufficiently fast. Details are found in Sec. $4.$
@@
Time series are an essential data modality ubiquitous in many critical fields of application, such as medicine (Hosseini et al., 2021), manufacturing (Essien and Giannetti, 2020), logistics (Seyedan and Mafakheri, 2020), traffic management (Lippi et al., 2013), finance (Lin et al., 2012; Divo et al., 2025), and weather modeling (Lam et al., 2023). While significant progress in time series forecasting has been made over the decades, the field is still far from being solved. Further increasing the forecast quality obtained from machine learning models promises a manifold of improvements, such as more accurate medical treatments, increased efficiency in manufacturing and transportation, and higher crop yields.
@@
Historically, recurrent neural networks (RNNs) and their powerful successors were natural choices for deep learning-based time series forecasting (Hochreiter and Schmidhuber, 1997; Cho et al., 2014). Today, large Transformers (Vaswani et al., 2017) are applied extensively to time series tasks, including forecasting. Many improvements to the vanilla architecture have since been proposed, including patching (Nie et al., 2023), decompositions (Zeng et al., 2023), and tokenization inversions (Liu et al., 2023). Newer approaches include pretrained models, which, however, usually cannot capture relationships between multiple variables (Ansari et al., 2024). Yet, some fundamental limitations of Transformers are yet to be lifted. For instance, they are inefficient when applied to long sequences due to the cost of the attention mechanism being quadratic in the number of variates and time steps. As embedded devices and edge computing platforms are gaining importance, the demand for lightweight forecasting models that balance accuracy with minimal memory and computational overhead grows. Therefore, recurrent and state space models (SSMs) (Patro and Agneeswaran, 2025) are experiencing a resurgence of interest in overcoming such limitations. Specifically, Beck et al. (2024) revisited recurrent models by borrowing insights gained from Transformers in many domains, specifically natural language processing. They propose Extended Long Short-Term Memory (xLSTM) models as alternatives to current sequence models.
@@
We propose xLSTM-Mixer [^1] Code available at https://github.com/mauricekraus/xlstm-mixer, a new state-of-the-art method for time series forecasting using recurrent deep learning methods, which strikes a balance by providing strong forecasting accuracy while remaining highly efficient, as Fig. 1 shows. Architecturally, we combine the highly expressive xLSTM architecture with carefully crafted time, variate, and multi-view mixing. These operations regularize the training and limit the number of model parameters by weight-sharing, effectively improving the learning of features necessary for accurate forecasting. xLSTM-Mixer initially computes a channel-independent linear forecast shared over the variates. It is then up-projected to a higher hidden dimension and subsequently refined by an xLSTM stack. It performs multi-view forecasting by producing a forecast from the original and reversed up-projected embedding. The powerful xLSTM cells thereby jointly mix time and variate information to capture complex patterns from the data. Both forecasts are eventually reconciled by a learned linear projection into the final prediction, called view mixing. An overview of our method is shown in Fig. 2.
@@
Figure 2: The xLSTM-Mixer architecture consists of three stages: (1) An initial NLinear forecast assuming channel independence and performing time mixing; (2) subsequent joint mixing, which mixes variate and time information through crucial applications of sLSTM blocks; and (3) view mixing, where the two latent forecast views are reconciled into a coherent final forecast.
@@
    We propose xLSTM-Mixer, a state-of-the-art method for time series forecasting, for the first time combining recurrent deep learning with a mixing architecture.
@@
    We extensively compare xLSTM-Mixer with existing methods for multivariate long-term time series forecasting and perform in-depth model analyses. The experiments demonstrate that xLSTM-Mixer consistently excels in a wide range of benchmarks.
Outline.   In the upcoming Sec. 2, we introduce preliminaries to allow us to motivate and explain xLSTM-Mixer in Sec. 3. We then comprehensively evaluate its effectiveness and inner workings in Sec. 4. We finally review related work in Sec. 5 and close with a conclusion and outlook in Sec. 6.
@@ 2 Background
After introducing the notation used throughout this work, we review xLSTM blocks and discuss whether leveraging channel mixing or their independence is beneficial in time series models.
Notation.   In multivariate time series forecasting, the model is presented with a time series $\bm{X}=\left(\bm{x}_{1},\dots,\bm{x}_{T}\right)\in\mathbb{R}^{V\times T}$ consisting of $T$ time steps with $V$ variates each. Given this context, the forecaster shall predict the future values $\bm{Y}=\left(\bm{x}_{T+1},\dots,\bm{x}_{T+H}\right)\in\mathbb{R}^{V\times H}$ up to a horizon $H$. A variate (sometimes called a channel) can be any scalar measurement, such as the occupancy of a road or the temperature in a power plant. The measurements are assumed to be carried out jointly, such that the $T+H$ time steps reflect a regularly sampled signal. A time series dataset consists of NN such pairs $\left\{\left(\bm{X}^{(i)},\bm{Y}^{(i)}\right)\right\}_{i\in\{1,\dots,N\}}$ divided into train, validation, and test portions.
@@ 2.1 Extended Long Short-Term Memory (xLSTM)
Beck et al. (2024) propose xLSTM consisting of two building blocks, namely the sLSTM and mLSTM modules. To harness the full expressivity of xLSTMs within each step and across the computation sequence, we employ a stack of sLSTM blocks without any mLSTM blocks. The latter are less suited for joint mixing due to their independent treatment of the sequence elements, making it impossible to learn any relationships between them directly. See App. B for a deeper discussion. We will continue by recalling how sLSTM cells function. The standard LSTM architecture of Hochreiter and Schmidhuber (1997) involves updating the cell state $\mathbf{c}_{t}$ through a combination of input, forget, and output gates regulating the flow of information across tokens. sLSTM blocks enhance this by incorporating exponential gating and memory mixing (Greff et al., 2017) to handle complex temporal and cross-variate dependencies better. The sLSTM updates the cell $\bm{c}_{t}$ and hidden state $\bm{h}_{t}$ as follows:
$\bm{c}_{t}$ $=\bm{f}_{t}\odot\bm{c}_{t-1}+\bm{i}_{t}\odot\bm{z}_{t}$ cell state (1)
@@
$\bm{z}_{t}$ $=\tanh\bigl(\bm{W}_{z}\bm{x}_{t}+\bm{R}_{z}\bm{h}_{t-1}+\bm{b}_{z}\bigr)$ cell input (4)
@@
In this setup, the matrices $\bm{W}_{z},\bm{W}_{i},\bm{W}_{f},$ and $\bm{W}_{o}$ are input weights mapping the input token $\bm{x}_{t}$ to the cell input $\bm{z}_{t}$, input gate, forget gate, and output gate, respectively. The states $\bm{n}_{t}$ and $\bm{m}_{t}$ serve as necessary normalization and training stabilization, respectively. As Beck et al. have shown, it is beneficial to restrict the memory mixing performed by the recurrent weight matrices $\bm{R}_{z},\bm{R}_{i},\bm{R}_{f},$ and $\bm{R}_{o}$ to individual heads, inspired by the multi-head setup of Transformers (Zeng et al., 2023), yet more restricted and efficient. In particular, each token gets broken up into groups of features, where the input weights $\bm{W}_{z,i,f,o}$ act across all of them, but the recurrence matrices $\bm{R}_{z,i,f,o}$ are implemented as block-diagonal. This permits specialization of the individual heads to patterns specific to the respective section of the tokens and empirically does not sacrifice expressivity.
2.2 Channel Independence and Mixing for Time Series
Multiple works have investigated whether it is beneficial to learn representations of the time and variate dimensions jointly or separately. Intuitively, because joint mixing is strictly more expressive, one might think it should always be preferred. And, indeed, it is used in many works including Temporal Convolutional Networks (Lea et al., 2016), N-BEATS (Oreshkin et al., 2019), N-HiTS (Challu et al., 2023), and many Transformers (Lim et al., 2021; Wu et al., 2021; Zhou et al., 2022). However, treating slices of the input data independently assumes an invariance to temporal or variate positions and serves as a strong regularization against overfitting, reminiscent of kernels in CNNs. Prominent models implementing some aspects of channel independence in multivariate time series forecasting are PatchTST (Nie et al., 2023) and iTransformer (Liu et al., 2023). TiDE (Das et al., 2023), on the other hand, contains a time-step shared feature projection and temporal decoder but treats variates jointly. As Tolstikhin et al. (2021) have shown with MLP-Mixer, interleaving mixing of all channels per token and all tokens per channel does not empirically sacrifice any expressivity and instead improves performance and efficiency. This idea has since been applied to time series too, namely in architectures such as TimeMixer(++) (Wang et al., 2024a, 2025a), TSMixer (Chen et al., 2023c), WPMixer (Murad et al., 2025), and PMformer (Lee et al., 2024), and is, therefore, one key component of our method xLSTM-Mixer.
@@ 2.2 Channel Independence and Mixing for Time Series
We now explain xLSTM-Mixer shown in Fig. 2 in more detail. It carefully integrates several key components: an initial linear forecast with time mixing, joint mixing using powerful sLSTM blocks, and an eventual combination of two views by a final fully connected layer. The transposing steps between the components enable capturing complex temporal and intra-variate patterns while facilitating easy trainability and limiting parameter counts. The sLSTM blocks, in particular, can learn intricate non-linear relationships hidden within the data along both the time and variate dimensions. The architecture is furthermore equipped with normalization layers and skip connections to improve training stability and overall effectiveness.
3.1 Normalization and Initial Linear Forecast
Normalization has become an essential ingredient of modern deep learning architectures (Huang et al., 2023). For time series in particular, reversible instance norm (RevIN) (Kim et al., 2022) is a general recipe for improving forecasting performance, where each time series instance is normalized by its mean and variance and furthermore scaled and offset by a learnable scale $\bm{\gamma}$ and offset $\bm{\beta}$:
$\bm{x}_{t}^{\text{norm}}=\operatorname{RevIN}(\bm{x}_{t})=\bm{\gamma}\odot\left(\frac{\bm{x}_{t}-\mathbb{E}\left[\bm{x}\right]}{\sqrt{\mathrm{Var}\left[\bm{x}\right]}+\bm{\epsilon}}\right)+\bm{\beta}.$
We apply it as part of xLSTM-Mixer, and at the end of the entire pipeline, we invert the RevIN operation to obtain the final prediction. In the case of xLSTM-Mixer, the typical skip connections found in mixer acrchitectures (Tolstikhin et al., 2021; Chen et al., 2023c) are taken up by RevIN, the normalization in the NLinear forecast (Zeng et al., 2023), and the integral skip connections within each sLSTM block.
@@ 3.1 Normalization and Initial Linear Forecast
where $\operatorname{FC}(\cdot):\mathbb{R}^{T}\to\mathbb{R}^{H}$ denotes a fully-connected linear layer with bias term. Sharing this model across variates limits parameter counts, and the weight-tying serves as a useful regularization. The quality of this initial forecast will be investigated in Sec. 4.1 and 4.5.
3.2 sLSTM Refinement
@@
While the NLinear forecast $\bm{x}^{\text{initial}}\in\mathbb{R}^{V\times H}$ captures the basic patterns between the historic and future time steps, its quality alone is insufficient for today’s challenging datasets. We, therefore, refine it using powerful sLSTM blocks. As a first step, it is crucial to increase the embedding dimension of the data to provide sufficient latent dimensions $D$ for the sLSTM cells as $\bm{x}^{\text{up}}=\operatorname{FC}^{\text{up}}\left(\bm{x}^{\text{initial}}\right)$. This prior up-projection is similar to what is commonly performed in SSMs (Beck et al., 2024). We weight-share $\operatorname{FC}^{\text{up}}:\mathbb{R}^{H}\to\mathbb{R}^{D}$ across variates to perform time-mixing similar to the initial forecast. Note that this step does not maintain the temporal ordering within the embedding token dimensions, as was the case up until this step, and instead embeds it into a higher latent dimension.
@@
The stack of $M$ sLSTM blocks $\mathcal{S}(\cdot)$ transforms $\bm{x}^{\text{up}}\in\mathbb{R}^{V\times D}$ as defined in Eq. 1 to 8. The recurrent model strides over the data in variate order, i.e., where each token represents all time steps from a single variate as in the work of Liu et al. (2023). The sLSTM blocks learn intricate non-linear relationships hidden within the data along both the time and variate dimensions. The mixing of the hidden state is still limited to blocks of consecutive dimensions, aiding efficient learning and inference while allowing for effective cross-variate interaction during the recurrent processing. Striding over variates has the benefit of linear time scaling in the number of variates at a constant number of parameters. It, however, comes at the cost of possibly fixing a suboptimal order of variates. While this is empirically not a significant limitation (see Sec. 4.4), we leave investigations into how to find a suitable ordering for future work. In addition to a large embedding dim, we observed a high number of heads being crucial for effective forecasting.
@@
The sLSTM cells’ first hidden state $\bm{h}_{0}$ must be initialized before each sequence of tokens can be processed. Extending the initial description of these blocks, we propose learning a single initial embedding token $\bm{\eta}\in\mathbb{R}^{D}$ that gets prepended to each encoded time series $\bm{x}^{\text{up}}$. These initial embeddings draw from recent advances in Large Language Models, where learnable "soft prompt" tokens are used to condition models and improve their ability to generate coherent outputs (Lester et al., 2021; Li and Liang, 2021; Chen et al., 2023a, b). Recent research has extended the application of soft prompts to LLM-based time series forecasting (Cao et al., 2023; Sun et al., 2024; Jin et al., 2024), emphasizing their adaptability and effectiveness in improving model performance across modalities. These tokens enable greater flexibility by conditioning its initial memory representation to specific dataset characteristics for dynamically interacting with the time and variate data. Soft prompts can be readily optimized through back-propagation with very little overhead.
3.3 Multi-View Mixing